OpenStudy (anonymous):

As a fundraiser, a school club is selling posters. The printer charges a $21 set up fee, plus $0.40 for each poster. Then the cost y in dollars to print x posters is given by the linear equation y=0.40x+21. a) What is the cost y in dollars to print 50 posters? To print 100 Posters? b) Find the number of posters x if the printer billed the club for costs of $81.00 c) Write the information from parts (a) and (b) as three ordered pairs. The first paid from (a) is The second pair from (a) is The paid from (b) is

OpenStudy (anonymous):

(a) $41 explanation: .4(50)+21=41 (a)$61 explanation: .4(100)+21=61 (c) 150 posters explanation: 150(.4)+21=$81

OpenStudy (kymber):

 For A: substitute 50 and 100 into the equation as x Solve For B: substitute $81 dollars as y in the equation, since y is the total fee. Solve